Give the slope and y-intercept of the line then graph. y = 4x -3
Slope
y-intercept
x=3 graph on graph paper
y= -4 graph on graph paper

Answer by ewatrrr(24785)


Hi,
the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept. y = 4x -3 slope= 4
green line is y = 4x-3, x = 3 is the blue line and y =-4 is the red line